OpenStudy (anonymous):

-7(t-4)- (t+9) =5 distributive I got t=4 ??

OpenStudy (whpalmer4):

Does t = 4 work in the original equation? \[-7(4-4)-(4+9)=5\]\[-7(0)-(13)=5\]\[-13=5\]Not in my universe :-)

OpenStudy (whpalmer4):

\[-7(t-4)-(t+9)=5\] \[-7t+28 -t -9 = 5\]Note that -7(-4) = 28 and -(+9) = -9 — frequent source of error! Now finish solving for t, what do you get?

OpenStudy (anonymous):

then I get 8t+37=5

OpenStudy (whpalmer4):

Check your signs again...

OpenStudy (whpalmer4):

Pun intended :-)

OpenStudy (anonymous):

-8t+37=5?

OpenStudy (whpalmer4):

\[-7t+28-t-9 = -7t-t+28-9 = -8t + 19\] right?

OpenStudy (anonymous):

ok so minus instead of adding the 9 to each side

OpenStudy (anonymous):

so -2

OpenStudy (whpalmer4):

You can add 9 to both sides:\[-7t+28 -t -9 +9 = 5 + 9\] And subtract 28 from both sides:\[-7t-t+28=14\]\[-8t+28-28=14-28=-14\]\[-8t=-14\]

OpenStudy (whpalmer4):

What does t equal? Final answer :-)

OpenStudy (anonymous):

1 3/4 ?

OpenStudy (whpalmer4):

Once again, let's check our solution: \[-7(t-4)- (t+9) =5\]\[-7(1\frac{3}{4}-4)-(1\frac{3}{4}+9)=5\]\[-7(-2\frac{1}{4})-10\frac{3}{4}=5\]\[14\frac{7}{4}-10\frac{3}{4}=5\]\[4\frac{4}{4}=4+1=5\] Looks like that is correct, good job!